F. None of the above Question 18 Your favorite model of car sells now for 350,000 SEK and this price will increase by the annual inflation year after year. You would like to purchase it when you get your driving license in 3 years. lo make sure that you dont spend the money in your wild partying during those three years, you decide to make a one-off deposit of money today in a savings account whose conditions do not allow any withdrawal over the entire three-year period. Suppose that you will be taxed on your interest at a rate equal to 25% and that the nominal interest rate that you will be earning on your deposit is equal to 4% (EAR). If the inflation rate per year is equal to 2%, what will be the money that you need to deposit today so that you can have in your savings account exactly the amount needed to purchase the car in three years? (Answers rounded to a whole number)

hello i got the answer but not tye solutions so please dont try to bw funny. so can anyone explain how to solve this step by step?
Answer #1
Ampount to be accumulated considering inflation = 350000*1.02^3 = $   3,71,423
Interest rate earned net of tax = 4%*(1-25%) = 3.00%
Amount to be deposited today = 371423/1.03^3 = $   3,39,905